OpenStudy (anonymous):

Which is the area of a circle with a circumference of 15.7 units? (use 3.14 for pie)

OpenStudy (callisto):

Hmm.. It is *pi* not *pie*!!!!!!! :'( Circumference = 2(pi)r, where r is the radius, pi is 3.14 Area = (pi)(r)(r) Since you know the circumference is 15.7, so 15.7 = 2(3.14) r Calculate r, and sub. it into the formula to calculate the area, that is Area = (pi)(r)(r)

OpenStudy (anonymous):

since the formula of circumference of a circle is 2(pi)r, so 2 (pi) r = 15.7 r = 15.7 / 2 (pi) r = 2.499 units substitute "r" to the formula for area of the circle A = (pi) r^2 = (pi) (2.499)^2 A = 19.61 sq.units

OpenStudy (aravindg):

Hi there! Welcome to Openstudy :) If you have any doubts on how the site works ,feel free to ask me or any other Ambassador with purple A near them .Have great time here . Check out our COC : http://openstudy.com/code-of-conduct

OpenStudy (aravindg):

@Callisto PIE :D

OpenStudy (callisto):

@pingube Please do not give the answer :( We are here to teach + learn